OpenStudy (thomasfire55):

If the length of one leg of a right triangle is 16 and the length of the other leg is odd, what are the lengths of all the sides of the triangle?

OpenStudy (prathamesh_m):

I think the only way you can find the other two sides if 16 is the hypotenuse.

OpenStudy (prathamesh_m):

Since u know that \[Hypotenuse ^{2}=a ^{2}+b ^{2}\] Find the value of 16 squared and then use trial and error to figure out the other two sides.

OpenStudy (thomasfire55):

coulnt you use sine and cosine

OpenStudy (prathamesh_m):

But you don't know any angles right?

OpenStudy (thomasfire55):

no i dont

OpenStudy (thomasfire55):

not besides the right angle

OpenStudy (prathamesh_m):

So I think would be best to use trial and error. You don't have enough data.

OpenStudy (prathamesh_m):

I think you can form only one Pythagorean triplet using the number 16.

OpenStudy (thomasfire55):

what is it?

OpenStudy (prathamesh_m):

That you'll have to figure out using trial and error.

OpenStudy (prathamesh_m):

You already know that one member of the triplet is odd so that may be useful.

OpenStudy (thomasfire55):

16,63,65?

jhonyy9 (jhonyy9):

for example one pythagorean triplet is 3,4 and 5 and there are so much pythagorean triplets in the form of 3n,4n and 5n the given 16 is 4 squared hope this will help you deducting the right answer

OpenStudy (prathamesh_m):

Yeah @jhonyy9 is right

jhonyy9 (jhonyy9):

for example than you multiplie these 3,4,5 triplet members by 4 you get 3*4=12 4*4=16 4*5=20 hope these will help you sure easy

jhonyy9 (jhonyy9):

ty. @Prathamesh_M

OpenStudy (welshfella):

only you haven't got an odd length for the other leg.

OpenStudy (welshfella):

there maybe many solutions to this problem - the question doesn't say that all lengths are integers.

OpenStudy (welshfella):

You are correct thomas 65,63 and 16 fits the bill!

OpenStudy (welshfella):

- the question doesn't say that the hypotenuse cannot be odd.

OpenStudy (welshfella):

http://www.tsm-resources.com/alists/trip.html - that is a list of pythagorean triples